Ho, Lori & Sophia had 39 CDS among them. Ho gave 6 CDs to Lori & Sophia gave 2 CDs to Ho. Finally, Lori gave 3 CDs to Ho. At this point, they each had the same # of CDs. How many CDs did each person originally have?
Answer by Theo(13342)

when they are done trading CD's, they each have an equal number of CD's.
since the total number of CD's is equal to 39, this means that they each have 13 CD's.
Let H = number of CD's that Ho has when they started trading.
Let L = number of CD's that Lori has when they started trading.
Let S = number of CD's that Sophia has when they started trading.
you need to tabulate all the changes from when they started to get their starting numbers.
For Ho:
H - 6 + 2 + 3 = H - 1
H - 1 = 13 means that H = 14 when they started.
For Lori:
L + 6 - 3 = L + 3
L + 3 = 13 means that L = 10 when they started.
For Sophia:
S - 2 = S - 2
S - 2 = 13 means that S = 15 when they started.
you get:
Ho had 14 when they started.
Sophia had 15 when they started.
Lori has 10 when they started.
This is summarized as:
Ho 14 Sophia 15 Lori 10
Ho gave 6 CDs to Lori.
Tally is now:
Ho 8 Sophia 15 Lori 16
Sophia gave 2 CD's to Ho
Tally is now:
Ho 10 Sophia 13 Lori 16
Lori gave 3 CD's to Ho
Tally is now:
Ho 13 Sophia 13 Lori 13
They're done and they all have the same number of CD's.